Prediction Questions: The predictions provided on each lab handout are to help you prepare for the lab session. The questions will typically ask you to either do a calculation or think about a concept that will be used in the lab.

Objective:  The objective of this laboratory exercise is to understand the effect that gravity has on falling objects.   You will also see two ways of measuring the acceleration due to gravity.

Predictions:  A rock is dropped off a bridge and falls to the river below.  The rock falls a distance of 100 meters to the water.  Answer the following prediction questions and turn them in at the beginning of lab.

1) Draw graphs of vertical position, velocity and acceleration for the rock which include correct numerical axes for the motion.

2) Draw the graphs again if the rock is thrown downward at 10 m/s instead of dropped.
Answer #1

The equation for position, velocity and acceleration of the rock of a free fall (initial velocity is zero ) is

\Delta y = v_i\cdot t +\frac{1}{2}g t^2

100 = 0 +\frac{1}{2}(9.8) t^2

t = 4.5175s

The rock reaches the water in 4.5175s. Now using this as parameter we determine the graph

time (t)Position (y) (0.5)*(g)*t^2 velocity (v)-g tAcceleration (g) 120 0 0 4.9 19.6 44.1 78.4 99.99825063 0 9.8 19.6 29.4 39

b) The same is done with initial speed of 10m/s then time would be

\Delta y = v_i\cdot t +\frac{1}{2}g t^2

100 = 10 t +\frac{1}{2}(9.8) t^2

4.9 t^2 + 10 t - 100 = 0

solving for t we have

t = 3.61s